OpenStudy (anonymous):

Find two sets of coordinates for two lines that when drawn create a pair of lines. Show the lines are either parallel or not parallel by showing your work with the slope formula for each line. Then, write an equation of a line parallel to one of the lines in slope-intercept form that contains the point (1,-3). You must show all work and steps to receive credit.

OpenStudy (anonymous):

I am just confused how to write an equation of a line parallel to one of the lines in the slope-intercept form that contains the point (1, -3).

OpenStudy (anonymous):

@jagr2713

OpenStudy (anonymous):

Have you got any equation yet?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok, two parallel line must have the same slope . Let use a random one like Y=ax+b Let's use slope equals to 3 So Y=3x+b Substituting the point (1,-3) value we get this equation Y= 3x+10 A parallel equation would be Y=3x+4

OpenStudy (anonymous):

ok so I should use that as my equation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok, I have one line at (5, 36) and (12, 30), so what would be an equation that is parallel?

OpenStudy (anonymous):

The slope that i got for that line was 6/-7.

OpenStudy (anonymous):

Y=ax+b First point 36=5a+b Second point 30=12a+b System 36=5a+b 30=12a+b (-1) ------------------- 36=5a+b -30=-12a-b ---------------- 6=-7a a=-6/7 So 36=5.(-6/7)+b 36=-30/7+b b=252+30/7 b=282/7 Function y=-6/7x+282/7

OpenStudy (anonymous):

ok so would y=-6/7x-15/7 be the final answer?